The chemical formula for carbon dioxide is CO2. If the weight of carbon dioxide is 3520 grams, how much does the oxygen weigh? (C,12:O,16)
Answer by edjones(7311)

CO2=12+2(16)
=12+32
=44
.
Let x be the wt of O2
x/32=3520/44 Proportion
x=32*3520/44
x=2560 gms.
.
Ed